Planning week 21-28 May 2011

This week lies within the eclipse season.

Although there were several ARs on the disk this week, they were all very quiet. AR 11216 and 11218 passed the central meridian in the first half of the week and were the main focus of observations. At the end of the week we switched to another, unnumbered, active region at the same latitude that was approaching the central meridian.

There were only two HOPs in the plan: HOP 186 was related to the XRT synoptic and required no EIS support; HOP 81 was supported on 27 May by observing the north pole.

AR 11218 had a small coronal hole on its north-east side and fan loops could be seen extending over the coronal hole from the active region. I ran several studies over this region.

The week was divided into four OP periods: OP1 was 3 days long, OP2 1 day, OP3 2 days and OP4 1 day. 

Saturday 21 May (OP 1)

AIA 193 image at 23:44 UT

AR 11214 was close to the west limb and, although Hinode was tracking it, I performed a sit-and-stare above the limb with PRY_sit_stare_1 for 4 hours between 11:30 and 16:15.

Sunday 22 May (OP 1)

AIA 193 image at 23:42 UT

Pointing moved to AR 11216 (near disk center) and I ran a set of five IUU_SCAN_STEPS_002 sequences between 00:45 and 08:20 centered on the core of the active region. I then ran a single HPW021_VEL_120x512v2 raster at 08:56, and a ATLAS_60 at 10:33.

Pointing moved to disk center where I performed the sensitivity monitoring study (SYNOP006) at 12:12. The north and east co-alignment study was then performed during 15:55-19:40.

Monday 23 May (OP 1)

AIA 193 image at 23:44 UT

Pointing switched to AR 11218 today. Only one study was run: the hunter_loop_fe8 event trigger at 12:47.

Tuesday 24 May (OP 1 & OP 2)

AIA 193 image at 20:39 UT

Pointing continued on AR 11218. The Cambridge AR brightening study was run between 01:55 and 06:20 (3 orbits).

Between 10:30 and 16:00 (4 orbits) I ran Jim and Spiros' nanoflare study on AR 11218. The first two pointings were in the AR core; the second two were in the more peripheral loops (seen in AIA 335).

Another run of the Cambridge AR brightening study was scheduled at 16:42 (one orbit).

The HPW021_VEL_120x512v2 raster was run again at 23:28 and 01:03 (Wednesday). The two rasters were placed side-by-side and so should give a good map of the combined AR fan loops and coronal hole.

Wednesday 25 May (OP 2 & OP 3)

AIA 193 image at 23:37 UT

Pointing switched back to AR 11216 at SOT's request as there is a small sunspot. I took the opportunity to focus on the loops coming from the sunspot. I first ran EL_WHI_CH_limb which takes 3 exposures of 600 s (!) at a fixed position (11:00 UT). I then ran four consecutive Large_CH_Maps between 12:25 and 18:30 which should cover both the sunspot loops and some larger fan loops. I've scheduled the EL_WHI_CH_limb study again for the following orbit.

At 22:15 Hinode is moved to disk center (fixed pointing) for a SOT flat field. The coronal hole is quite close to disk center so I'm running EL_WHI_CH_limb as a "drift-and-stare" study for an hour which should yield a good coronal hole spectrum.

Thursday 26 May (OP 3)

AIA 193 image

The Hinode pointing moved to the coronal hole for a few hours, and I scheduled three consecutive Large_CH_Map rasters between 00:00 and 04:20. Note that these rasters will probably end up "contaminated" by the AR fan loops.

The co-alignment study was run again between 05:15 and 09:00.

Pointing moved back to AR 11218, and I scheduled three runs of Large_CH_Map on the AR fan loops (those extending north towards the equator) separated by a few hours. The first and second were at 11:25 and 19:38, and for each I used the following orbit to run EL_FULL_CCD_RASTER to obtain good reference spectra.

Friday 27 May (OP 3 & OP 4)

AIA 193 image

The third Large_CH_Map raster (see Thursday) was scheduled at 03:50.

HOP 81 (north pole) was supported with the SK_CH_DIAG_320x384 study at 11:55 (4.5 hour duration). I scheduled it so that it matched the period where SOT was taking SP scans. The period 10:42-11:30 was used to run PRY_slot_context in the same location.

Pointing switched from AR 11218 to an unnumbered AR at the same latitude that was approaching the central meridian. Just before the daily meeting it produced a few impulsive flares. The region has a hot core, with an interesting group of "converging" fan loops on the west side. I scheduled two runs of HPW021_VEL_120x512v2 to make a composite raster of the loops at 18:53 and 20:28.

Saturday 28 May (OP 4)

AIA 193 image

Three runs of nanoflare_sas_v1 were performed at 01:35, 03:08 and 04:43 in the core of the unnumbered active region. The first two pointings were at the footpoints of the brightest core loops, whereas the 3rd pointing was for the periphery loops (as seen in AIA 335).